# Exploring QUIC Dynamics: A Large-Scale Dataset for Encrypted Traffic Analysis

### v3

@@ Abstract.
QUIC, an increasingly adopted transport protocol, addresses limitations of TCP by offering improved security, performance, and features such as stream multiplexing and connection migration. However, these enhancements also introduce challenges for network operators in monitoring and analyzing web traffic, especially due to QUIC’s encryption. Existing datasets are inadequate—they are often outdated, lack diversity, anonymize critical information, or exclude essential features like SSL keys—limiting comprehensive research and development in this area. We introduce VisQUIC, a publicly available dataset of over 100,000 labeled QUIC traces with corresponding SSL keys, collected from more than 44,000 websites over four months. By generating visual representations of the traces, we facilitate advanced machine learning (ML) applications and in-depth analysis of encrypted QUIC traffic. To demonstrate the dataset’s potential, we estimate the number of HTTP/3 request-response pairs in a QUIC connection using only encrypted traffic, achieving up to 92% accuracy. This estimation provides insights into server behavior, client-server interactions, and connection load—crucial for tasks like load balancing and intrusion detection. Our dataset enables comprehensive studies on QUIC and HTTP/3 protocols and supports the development of tools for encrypted traffic analysis.
@@ 1. Introduction
The VisQUIC dataset results in over seven million labeled images, using two different window lengths, providing a rich resource for ML tasks in network analysis. By making this dataset publicly available, we facilitate comprehensive research on the behavior of HTTP/3 and QUIC, enabling advancements in network monitoring and security.
@@ 3.2. Image Generation Technique
youtube.com 399 2,109 139,889 54,659
semrush.com 1,785 9,489 474,716 221,477
discord.com 527 7,271 623,823 235,248
instagram.com 3 207 17,003 7,112
mercedes-benz.com 46 66 9,987 2,740
bleacherreport.com 1,798 8,497 781,915 331,530
nicelocal.com 1,744 1,666 148,254 48,900
facebook.com 13 672 25,919 10,988
pcmag.com 5,592 13,921 1,183,717 385,797
logitech.com 177 728 56,792 28,580
google.com 1,341 2,149 81,293 29,068
cdnetworks.com 902 2,275 207,604 85,707
independent.co.uk 3,340 3,453 176,768 68,480
cloudflare.com 26,738 44,700 1,347,766 341,488
jetbrains.com 35 1,096 34,934 18,470
pinterest.com 43 238 6,465 2,360
wiggle.com 4 0 0 0
cnn.com 27 2,127 91,321 59,671
@@ 4. Estimating the Number of HTTP/3 Responses in a QUIC Connection
Figure 3(b)(3(a)) shows the scatter plot for predictions from the model trained using $T=0.1$-second window images, while Figure 3(b)(3(b)) displays results for the $T=0.3$-second window. The test dataset includes $12,520$ traces with an average of $21.2$ images per trace for $T=0.1$ seconds, and $12,142$ traces with an average of $7.5$ images per trace for $T=0.3$ seconds.
@@
Figures 3(b)(3(a)) and 3(b)(3(b)) illustrate notable differences in predictive behavior between models trained and evaluated with $T=0.3$-second and $T=0.1$-second window sizes. Specifically, diagonal patterns appear in the predictions of the $T=0.3$-second model that are absent in the $T=0.1$-second model’s predictions. This phenomenon exists for several reasons. When using a $T=0.1$ subdivision, a high percentage of images have lower class values, and the model trained on this dataset is very accurate for low-value classes. In contrast, a $T=0.3$ subdivision yields images with higher class values, increasing the variance of the true labels, and both models perform worse for higher-value classes compared to lower-value ones. Additionally, any incorrect prediction by either model contributes to an increase in the cumulative predictions for the remainder of the trace, thereby elevating the overall predicted values. The $T=0.1$-second model’s box plot is presented in Appendix ..

### v4

@@ Abstract.
QUIC, an increasingly adopted transport protocol, addresses limitations of TCP by offering improved security, performance, and features such as stream multiplexing and connection migration. However, these enhancements also introduce challenges for network operators in monitoring and analyzing web traffic, especially due to QUIC’s encryption. Existing datasets are inadequate—they are often outdated, lack diversity, anonymize critical information, or exclude essential features like SSL keys—limiting comprehensive research and development in this area. We introduce VisQUIC, a publicly available dataset of over 100,000 labeled QUIC traces with corresponding SSL keys, collected from more than 40,000 websites over four months. By generating visual representations of the traces, we facilitate advanced machine learning (ML) applications and in-depth analysis of encrypted QUIC traffic. To demonstrate the dataset’s potential, we estimate the number of HTTP/3 request-response pairs in a QUIC connection using only encrypted traffic, achieving up to 92% accuracy. This estimation provides insights into server behavior, client-server interactions, and connection load—crucial for tasks like load balancing and intrusion detection. Our dataset enables comprehensive studies on QUIC and HTTP/3 protocols and supports the development of tools for encrypted traffic analysis.
@@ 1. Introduction
The VisQUIC dataset results in over seven million labeled images, using two different window lengths, providing a rich resource for ML tasks in network analysis. By making this dataset publicly available  [^1] At: https://github.com/robshahla/VisQUIC, we facilitate comprehensive research on the behavior of HTTP/3 and QUIC, enabling advancements in network monitoring and security.
@@ 3.2. Image Generation Technique
youtube 399 2,109 139,889 54,659
semrush 1,785 9,489 474,716 221,477
discord 527 7,271 623,823 235,248
instagram 3 207 17,003 7,112
mercedes-benz 46 66 9,987 2,740
bleacherreport 1,798 8,497 781,915 331,530
nicelocal 1,744 1,666 148,254 48,900
facebook 13 672 25,919 10,988
pcmag 5,592 13,921 1,183,717 385,797
logitech 177 728 56,792 28,580
google 1,341 2,149 81,293 29,068
cdnetworks 902 2,275 207,604 85,707
independent 3,340 3,453 176,768 68,480
cloudflare 26,738 44,700 1,347,766 341,488
jetbrains 35 1,096 34,934 18,470
pinterest 43 238 6,465 2,360
wiggle 4 0 0 0
cnn 27 2,127 91,321 59,671
@@ 4. Estimating the Number of HTTP/3 Responses in a QUIC Connection
Figure 4(b)(4(a)) shows the scatter plot for predictions from the model trained using $T=0.1$-second window images, while Figure 4(b)(4(b)) displays results for the $T=0.3$-second window. The test dataset includes $12,520$ traces with an average of $21.2$ images per trace for $T=0.1$ seconds, and $12,142$ traces with an average of $7.5$ images per trace for $T=0.3$ seconds.
@@
Figures 4(b)(4(a)) and 4(b)(4(b)) illustrate notable differences in predictive behavior between models trained and evaluated with $T=0.3$-second and $T=0.1$-second window sizes. Specifically, diagonal patterns appear in the predictions of the $T=0.3$-second model that are absent in the $T=0.1$-second model’s predictions. This phenomenon exists for several reasons. When using a $T=0.1$ subdivision, a high percentage of images have lower class values, and the model trained on this dataset is very accurate for low-value classes. In contrast, a $T=0.3$ subdivision yields images with higher class values, increasing the variance of the true labels, and both models perform worse for higher-value classes compared to lower-value ones. Additionally, any incorrect prediction by either model contributes to an increase in the cumulative predictions for the remainder of the trace, thereby elevating the overall predicted values. The $T=0.1$-second model’s box plot is presented in Appendix ..